A nuclear scientist claimed that he has accurately measured the energy and momentum of a
radiation (e.g. Alpha, Beta, Gamma-ray, Neutron etc.) using his most sensitive Geiger Muller
counter. Can you believe his claim is correct? If yes, give a proper justification for his claim and
if no, then also prove that he is wrong
1
Expert's answer
2020-07-14T08:51:49-0400

In Quantum Mechanics two quantities can be measured simultaneously, if their operators commute.

The operator for the radiation energy (kinetic) is the following:


"\hat{T} = -\dfrac{\hbar}{2m}\Delta"

where "\Delta = \dfrac{\partial^2}{\partial x^2} + \dfrac{\partial^2}{\partial y^2} + \dfrac{\partial^2}{\partial z^2}" is the Laplacian.

The operator for the momentum is the following:


"\hat{\mathbf{p}} = -i\hbar\nabla"

where "\nabla = \left( \dfrac{\partial}{\partial x}, \dfrac{\partial}{\partial y}, \dfrac{\partial}{\partial z}\right)" is the nabla operator.

These operators commute if the following equation is true:

"\hat{T}\hat{\mathbf{p}} =\hat{\mathbf{p}}\hat{T}"

As far as both operators consist of a partial derivatives, the question is whether or not we can change the order of the derivatives. According to the Clairaut's theorem, we can do it. Thus:


"\hat{T}\hat{\mathbf{p}} = -\dfrac{\hbar}{2m}\Delta(-i\hbar\nabla) = \dfrac{i\hbar^2}{2m}\Delta(\nabla) = \dfrac{i\hbar^2}{2m}\nabla(\Delta) = \hat{\mathbf{p}}\hat{T}"

Thus, operators comute and we can measure the energy and momentum of a radiation at the same time.


Answer. Yes, his claim is correct.